OpenStudy (anonymous):

Given cos t= 1/2, tan > 0, find sin (t) and sec (t). A. Sin (t)= - Square root3/2, sec (t) =2 B. Sin (t)= Square root3/2, sec (t) =-2 C. Sin (t)= - Square root3/2, sec (t) =-2 D. Sin (t)= Square root3/2, sec (t) =2

OpenStudy (danjs):

If the cosine is 1/2, then the sin will be... \[\sin^{2} x + \cos^{2}x = 1 \] fill in for the cos(x) = 1/2 and find the sin(x)

OpenStudy (danjs):

\[\sin^2(t) +\frac{ 1 }{ 4 } = 1 \]

OpenStudy (danjs):

the sec(t) is also 1 / cos(t)

OpenStudy (danjs):

Or another method, from a triangle 

OpenStudy (danjs):

Use pythagorean theorem to get the ? . 2^2 = 1^2 + ?^2 ? = then sin(t) = ? / 2

OpenStudy (danjs):

\[\sec(t) = \frac{ 1 }{ \cos(t) } = \frac{ 1 }{ 1/2 }\]

OpenStudy (anonymous):

This makes no sense to me

OpenStudy (danjs):

just use the relationship identity, \[\sin^2(t) + \cos^2(t) = 1\] cos(t) = 1/2 Given \[\sin^2(t) + [\frac{ 1 }{ 2 }]^2 = 1\] \[\sin^2(t) = 1 - \frac{ 1 }{ 4 }\] \[\sin^2(t) = \frac{ 3 }{ 4 }\] \[\sin(t) = \sqrt{\frac{ 3 }{ 4 }} = \frac{ \sqrt{3} }{ \sqrt{4} } = \frac{ \sqrt{3} }{ 2 }\] So the sin(t) is \[\sin(t) = \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (danjs):

For the secant of t, \[\sec(t) = \frac{ 1 }{ \cos(t) } = \frac{ 1 }{ 1/2 } = 2\] sec(t) = 2

OpenStudy (anonymous):

Ok know that makes more sense but how did that one become a 3?

OpenStudy (danjs):

That part? \[1 - \frac{ 1 }{ 4 } ~~=~~ \frac{ 4 }{ 4 } - \frac{ 1 }{ 4 } ~~= ~~\frac{ 3 }{ 4 }\]

OpenStudy (danjs):

u understand it? i am about to log out

OpenStudy (anonymous):

Yes. Thank you